40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 37     Entry time: Wed Oct 31 09:45:28 2007
Author: waldman 
Type: Other 
Category: OMC 
Subject: Resolution to DAQland saga 
[Jay, Sam]

We did a rough accounting for the linear delay this morning and it comes out more or less correct. The 10 kHz 3rd order butterworth AA/AI filter gives ~90 degrees of phase at 6 kHz, or 42 microseconds. Taken together, the two AA and AI filters are worth 80 microseconds. The 1.5 sample digital delay is worth 1.5/32768 = 45 microseconds. The remaining 160 - 125 = 35 microseconds is most likely taken up by the 64 kHz to 32 kHz decimation routine, assuming this isn't accounted for already in the 1.5 sample digital delay.

It remains to be seen whether this phase delay is good enough to lock the laser to the OMC cavity
ELOG V3.1.3-